42. Is –4 a solution to the equation 7x – 5 + 3x = 6 + x – 10
Thanks!!!!!!!!!!!

Found 2 solutions by checkley71, rcmcc:
Answer by checkley71(8403)

7X-5+3X=6+X-10
7*-4-5+3*-4=6-4-10
-28-5-12=6-14
-45=-8 THUS (-4) IS NOT A SOLUTION.

Answer by rcmcc(152)

7x – 5 + 3x = 6 + x – 10
is -4 the solution... No.
when you substitute in -4 for x you recieve the answer -35=-8, they are unequal.
7x – 5 + 3x = 6 + x – 10
first you can combine like terms
10x-5=x-4
remove the x from the right hand side and you get
9X-5=-4
remove the -5 from the right hand side
you get the equation
9x=1
divide by 9
x=(1/9)
when you substitute in to the equation (1/9) for x you get the statement
-3.8888=-3.8888